# Recurrent neural network analysis of single trajectories switching between anomalous diffusion states

### v1

@@ 4.3 Changepoint detection and location performances
Figure 4: Changepoint detection comparison: sequandi vs. CPDA Using the trajectory illustrated in Fig. 1, the two vertical solid black lines are sequandi’s prediction $\hat{t}^{\textrm{CP}}$. (a) As in Fig. 3, the orange line is $\hat{\alpha}(t)$, and the blue piece-wise constant line is the ground truth $\alpha(t)$. The three vertical dashed gray lines are obtained by the CPDA algorithm described in Section 4.4 applied to $\hat{\alpha}(t)$ with a confidence level of $99.99\%$. (b) As in Fig. 3, the orange line is $\log(\hat{K}(t)+1)$ and the blue piece-wise constant line is the ground truth $\log(K(t)+1)$. The six vertical dashed gray lines are obtained by the CPDA algorithm described in Section 4.4 applied to $\log(\hat{K}(t)+1)$ with a confidence level of $99.99\%$.
@@
An alternative way of detecting and locating a CP is to exploit the pointwise time-dependent prediction of $\hat{\alpha}(t)$ and $\hat{K}(t)$, and apply a changepoint detection algorithm, as done, for instance, in Ref. [56]. Here, we apply the Changepoint Detection Algorithm (CPDA) developed in Ref. [77] to the time series $\hat{\alpha}(t)$ and $\hat{K}(t)$ obtained in Section 3. The authors of Ref. [77] improved the method proposed in Ref. [87], casting CP detection as a hypothesis testing problem and applied it to analyse the rotary motion of the F1 molecular motor. The algorithm requires setting a confidence level for accepting the hypothesis that there was a changepoint, as we discuss in C. Setting the optimal confidence level requires balancing the trade-off between missing CPs (false negative) and wrongly detecting CPs when there was not one (false positive). The optimal choice depends on the data under consideration as we discuss in C and Fig.12. For the benchmark dataset, we find a confidence level of $99.99\%$ returns the best performance in terms of JSC. The CPDA applied to the temporal profile $\hat{\alpha}(t)$ returns consistent results, even though it displays a tendency to detect false positives CPs, as exemplified in Fig.4(a). The $\hat{K}(t)$ profile is less suitable for the CPDA as it accentuates the tendency to return false positives, as shown in Fig.4(b). We, therefore, focus on the performance of the CPDA applied to the $\hat{\alpha}(t)$ profile. The CDPA does not reach the performance of sequandi and achieves a $\textrm{JSC}_{\textrm{CP}}=0.2470$ (roughly half of the one of sequandi) and $\textrm{RMSE}_{\textrm{CP}}=3.065$ (roughly twice the one of sequandi). However, despite its sensitivity to fluctuations in $\hat{\alpha}(t)$ and its tendency to return false positive detections, the CPDA outperforms sequandi for trajectories displaying many CPs, as we discuss in Section 5.
@@ 5.4 Benchmark dataset
  • Experiment 1: A 3-state model inspired by the experiments of [88]. All states are characterised by $\alpha$ close to the normal diffusion $\alpha\sim 1$ but each state has a different $K$, see the $x$ axis of Fig. 13 and 14. Switching between states occurs rapidly.
@@
The distribution of the various diffusive parameters can be seen by observing the ground truth values of the plots in Figs. 13 and 14.
@@
Overall, for the inference of $\alpha$ and $K$, sequandi performs well and evenly throughout different types of experiments with two exceptions. Experiment 9 is the most challenging one because of its unique combination of immobile states and nearly ballistic ones that are edge cases in the training set, see Fig. 9. Experiment 8 is also challenging for the $K$ inference because it displays a single state but covers a vast range of $K$, which was not optimised during training (Compare Fig. 9 with Fig. 14).
In Figs. 13 and 14, we report how individual $\hat{\alpha}_{i}$ and $\hat{K}_{i}$ predictions compare to their ground truths on an experiment-wise basis.
Figure 8: sequandi and CPDA performances on CP detection and location on each experiment of the benchmark dataset (a) Comparison of Jaccard Similarity Coefficient of changepoint detections $\textrm{JSC}_{\textrm{CP}}$, between the CPDA with confidence level 99.99, and by the sequandi neural network. $0\leq\textrm{JSC}\leq 1$, with 1 being perfect detection. (b) Root Mean Squared Error of true positive changepoint detected $\textrm{RMSE}_{\textrm{CP}}$, by sequandi and the CPDA.
@@ 6 Conclusion and discussion
This article presents sequandi, a method based on LSTM recurrent neural networks to study single trajectories undergoing anomalous diffusion and switching between different diffusive states. Leveraging the properties of RNNs, sequandi can analyse trajectories of varying length and returns a point-wise time-dependent estimate of the diffusive properties of the trajectory in terms of its anomalous diffusion exponent $\alpha$ and generalised diffusion coefficient $K$, $\hat{\alpha}(t)$ and $\hat{K}(t)$, respectively. Such a time-dependent estimate is well-suited to study trajectories that irregularly switch between different diffusive states. We complement this analysis by training another network to detect and locate the changepoints when a switch between different diffusive states occurs $\hat{t}^{\textrm{CP}}$. We compare the performance of this neural network to that of the Change Point Detection Algorithm (CPDA) proposed in Ref. [77] applied to the time-dependent estimate $\hat{\alpha}(t)$. At variance with the network, the CPDA requires carefully setting a confidence level, which is a non-trivial step, whose optimal solution depends on the task at hand. In general, sequandi outperforms the CPDA applied to $\hat{\alpha}(t)$. A possible way to improve the CPDA performance is to apply it to a combination of $\hat{\alpha}(t)$ and $\hat{K}(t)$ or to use also the uncertainty uncertainty $\sigma^{2}_{\alpha}(t)$, which one would have to compute, as done in Ref. [55, 56].
We benchmark the performances of sequandi on the 2024 Anomalous Diffusion Challenge dataset, which is based on detecting the changepoints and predicting the $\alpha$ and $K$ for each segment between changepoints. This requires combining the results of the network that detects the changepoints $\hat{t}_{\textrm{CP}}$ with the one that infers $\hat{\alpha}(t)$, and $\hat{K}(t)$ and then taking an average over the identified segments. This procedure is not error-free and degrades the performances of the method.
@@ Acknowledgements
SB wishes to thank Chun-Biu Li for discussions on the changepoint algorithm of Ref. [77]. XQ was financially supported by the China Scholarship Council through the King’s-China Scholarship Council PhD Scholarship programme (K-CSC).
Appendix A Training Dataset
In this Appendix, we provide additional details about the structure of the $2.5\times 10^{6}$ trajectories that we used for training sequandi. To generate the training set, we first uniformly sampled $\alpha$ and $\log{K}$ and generated the trajectories associated with these values. We then added trajectories with $\alpha$ close to 0 and 1 and small $K$ to have more training data for the most challenging cases. The resulting distributions are shown in Fig. 9 together with the corresponding segment-wise predictions
@@ Appendix A Training Dataset
Figure 9: Training trajectory dataset distributions and predictions. (a) Predicted segment-wise anomalous diffusion exponent $\hat{\alpha}_{i}$ vs the ground truth $\alpha_{i}$ for the training dataset of sequandi. (b) Predicted segment-wise generalised diffusion coefficient $\hat{K}_{i}$ vs the ground truth $K_{i}$. The bottom row of histograms show the distribution of the ground truth parameters of the dataset. Note the logarithmic scaling of $K$.
Appendix B Neural network changepoint detection signal
@@
Appendix C Changepoint algorithm optimisation
@@
    The first step checks if there are changepoints in a trajectory segment. Here, we are using as input the prediction from our networks $\hat{\alpha}(t)$. A straight line $\alpha_{\text{fit}}(t)$ is fitted. The residual is computed as $d(t)=\hat{\alpha}(t)-\alpha_{\text{fit}}(t)$, followed by the cumulative sum $\text{CUMSUM}(t)$ of $d(t)$. The statistic $D_{\text{data}}$ is the difference between the maximum and minimum values of $\text{CUMSUM}(t)$. The significance of $D_{\text{data}}$ is tested by permuting $d(t)$ multiple times to generate a distribution of $\text{CUMSUM}(t)$ curves and $D_{\text{permute}}$ values. A confidence level (e.g., 95%) sets the threshold $D_{\text{cutoff}}$. If $D_{\text{data}}\geq D_{\text{cutoff}}$, a changepoint is present. We discuss the relevance of the confidence level in C.1
@@
C.1 Confidence levels for the CPDA

### v2

@@ 4.3 Changepoint detection and location performances
Figure 4: Changepoint detection comparison: sequandi vs. CPDA Using the trajectory illustrated in Fig. 1, the two vertical solid black lines are sequandi’s prediction $\hat{t}^{\textrm{CP}}$. (a) As in Fig. 3, the orange line is $\hat{\alpha}(t)$, and the blue piece-wise constant line is the ground truth $\alpha(t)$. The three vertical dashed gray lines are obtained by the CPDA algorithm described in Section 4.4 applied to $\hat{\alpha}(t)$ with a confidence level of $99\%9$. (b) As in Fig. 3, the orange line is $\log(\hat{K}(t)+1)$ and the blue piece-wise constant line is the ground truth $\log(K(t)+1)$. The six vertical dashed gray lines are obtained by the CPDA algorithm described in Section 4.4 applied to $\log(\hat{K}(t)+1)$ with a confidence level of $99\%9$.
@@
An alternative way of detecting and locating a CP is to exploit the pointwise time-dependent prediction of $\hat{\alpha}(t)$ and $\hat{K}(t)$, and apply a changepoint detection algorithm, as done, for instance, in Ref. [61]. Here, we apply the Changepoint Detection Algorithm (CPDA) developed in Ref. [88] to the time series $\hat{\alpha}(t)$ and $\hat{K}(t)$ obtained in Section 3. The authors of Ref. [88] improved the method proposed in Ref. [98], casting CP detection as a hypothesis testing problem and applied it to analyse the rotary motion of the F1 molecular motor. The algorithm requires setting a confidence level for accepting the hypothesis that there was a changepoint, as we discuss in D. Setting the optimal confidence level requires balancing the trade-off between missing CPs (false negative) and wrongly detecting CPs when there was not one (false positive). The optimal choice depends on the data under consideration as we discuss in D and Fig.14. For the benchmark dataset, we find a confidence level of $99\%9$ returns the best performance in terms of JSC. The CPDA applied to the temporal profile $\hat{\alpha}(t)$ returns consistent results, even though it displays a tendency to detect false positives CPs, as exemplified in Fig.4(a). The $\hat{K}(t)$ profile is less suitable for the CPDA as it accentuates the tendency to return false positives, as shown in Fig.4(b). We, therefore, focus on the performance of the CPDA applied to the $\hat{\alpha}(t)$ profile. The CDPA does not reach the performance of sequandi and achieves a $\textrm{JSC}_{\textrm{CP}}=0.2470$ (roughly half of the one of sequandi) and $\textrm{RMSE}_{\textrm{CP}}=3.065$ (roughly twice the one of sequandi). However, despite its sensitivity to fluctuations in $\hat{\alpha}(t)$ and its tendency to return false positive detections, the CPDA outperforms sequandi for trajectories displaying many CPs, as we discuss in Section 5.
@@ 5.4 Benchmark dataset
  • Experiment 1: A 3-state model inspired by the experiments of [99]. All states are characterised by $\alpha$ close to the normal diffusion $\alpha\sim 1$ but each state has a different $K$, see the $x$ axis of Fig. 17 and 18. Switching between states occurs rapidly.
@@
The distribution of the various diffusive parameters can be seen by observing the ground truth values of the plots in Figs. 17 and 18.
@@
Overall, for the inference of $\alpha$ and $K$, sequandi performs well and evenly throughout different types of experiments with two exceptions. Experiment 9 is the most challenging one because of its unique combination of immobile states and nearly ballistic ones that are edge cases in the training set, see Fig. 9. Experiment 8 is also challenging for the $K$ inference because it displays a single state but covers a vast range of $K$, which was not optimised during training (Compare Fig. 9 with Fig. 18). In Figs. 17 and 18, we report how individual $\hat{\alpha}_{i}$ and $\hat{K}_{i}$ predictions compare to their ground truths on an experiment-wise basis.
[…]
@@ 6 Conclusion and discussion
This article presents sequandi, a method based on LSTM recurrent neural networks to study single trajectories undergoing anomalous diffusion and switching between different diffusive states. Leveraging the properties of RNNs, sequandi can analyse trajectories of varying length and returns a point-wise time-dependent estimate of the diffusive properties of the trajectory in terms of its anomalous diffusion exponent $\alpha$ and generalised diffusion coefficient $K$, $\hat{\alpha}(t)$ and $\hat{K}(t)$, respectively. Such a time-dependent estimate is well-suited to study trajectories that irregularly switch between different diffusive states. We complement this analysis by training another network to detect and locate the changepoints when a switch between different diffusive states occurs $\hat{t}^{\textrm{CP}}$. We compare the performance of this neural network to that of the Change Point Detection Algorithm (CPDA) proposed in Ref. [88] applied to the time-dependent estimate $\hat{\alpha}(t)$. At variance with the network, the CPDA requires carefully setting a confidence level, which is a non-trivial step, whose optimal solution depends on the task at hand. In general, sequandi outperforms the CPDA applied to $\hat{\alpha}(t)$. A possible way to improve the CPDA performance is to apply it to a combination of $\hat{\alpha}(t)$ and $\hat{K}(t)$ or to use also the uncertainty uncertainty $\sigma^{2}_{\alpha}(t)$, which one would have to compute, as done in Ref. [60, 61]. Another possibility would be pooling the CP detections of sequandi and the CPDA. This would require careful tuning, but it has the potential to mitigate the opposite tendencies towards false positives and false negatives the two approaches show in Fig. 8.
We benchmark the performances of sequandi on the 2024 Anomalous Diffusion Challenge dataset, which is based on detecting the changepoints and predicting the $\alpha$ and $K$ for each segment between changepoints. This requires combining the results of the network that detects the changepoints $\hat{t}_{\textrm{CP}}$ with the one that infers $\hat{\alpha}(t)$, and $\hat{K}(t)$ and then taking an average over the identified segments. This procedure is not error-free and degrades the performances of the method. However, sequandi was significantly outperformed only by the method of Ref. [85] and performed comparably to the runner-up methods (see Fig. 6).
@@
Since some experimental data comes from three-dimensional tracking, it may be interesting to extend the method to address three-dimensional trajectories. This extension would simply require adapting the data pre-processing and training the networks on three-dimensional data. It would also be interesting to explore the performance of different neural network architectures. Among these, transformers [101] may be especially promising, given their success in sequence modelling tasks, including natural language processing and time series forecasting. Their application to the study of anomalous diffusion has not been straightforward [77] due to difficulties in encoding the input trajectories [68], which have been mitigated by hybrid approaches, including initial CNN layers [79, 72, 86]. The AnDi2024 database will provide an objective testbed to compare the capabilities of different approaches, and has already highlighted the potential of CNN [85].
@@ Acknowledgements
SB wishes to thank Chun-Biu Li for discussions on the changepoint algorithm of Ref. [88]. XQ was financially supported by the China Scholarship Council through the King’s-China Scholarship Council PhD Scholarship programme (K-CSC). This work benefited from access to the Computational Research, Engineering and Technology Environment (CREATE) at King’s College London. We are grateful to the UK Materials and Molecular Modelling Hub for computational resources, which is partially funded by EPSRC (EP/T022213/1, EP/W032260/1 and EP/P020194/1. We acknowledge the support of Alejandro Santana Bonilla.
Appendix A Training and Testing Procedures
In this Appendix, we provide additional details about the structure of the $2.5\times 10^{6}$ trajectories that we used for training sequandi. and, analysing an additional, independent test dataset, we show that the network is correctly trained (not overfitting).
@@ Appendix A Training and Testing Procedures
Figure 9: Training dataset distributions and predictions. Bottom row histograms: distribution of the ground truth point-wise parameters of sequandi’s training dataset, (a) anomalous diffusion exponent $\alpha(t)$ and (b) generalised diffusion coefficient $K(t)$. Note the logarithmic scaling of $K$. Density maps: Predicted $\hat{\alpha}(t)$ and $\hat{K}(t)$ vs their respective ground truths. Colours are normalised in each column with respect to the ground truth probabilities below. Top row histograms: average mean absolute error (MAE) and mean squared logarithmic error (MSLE), of the points in each ground truth column. Per trajectory, the error averages over the entire dataset are $\textrm{MAE}_{\alpha(t)}=0.247$ and $\textrm{MSLE}_{K(t)}=0.140$.
To generate the training set, we first uniformly sample from the five experiments in the andi-datasets Python library in $\alpha$ and $\log{K}$. As mentioned in the main text, in supervised regression tasks, the values at the edges of the considered intervals are often difficult to estimate and an “S-shape” often occurs, where the lowest values are overestimated and the highest ones underestimated [57]. To mitigate this issue and to have more training data for the most challenging cases, which where frequent in the Andi2024 dataset, we add trajectories with $\alpha$ close to 2 and small $K$. The parameter distributions are shown in Fig. 9 together with the corresponding point-wise predictions. From the distributions in Fig. 9(a), we see the network excels for the central part of the $\alpha$ interval but underestimates the near ballistic case with $\alpha\approx 2$. This is confirmed by the $\alpha$ dependence of the $\textrm{MAE}_{\alpha(t)}$, (the upper panel of Fig. 9(a)), which shows how the network performs best in the middle of the $\alpha$ range. For the generalised diffusion coefficient $K$, the network trained well for $K\lesssim 10$ but systematically underestimates for $K>10$, as shown in Fig. 9(b). This is likely because, despite sampling evenly the number of segments with $K_{i}$ in this regime, trajectories with high $K$ are shorter, because they leave the simulation’s field of view more quickly. As a consequence, the piece-wise trajectories $K(t)$ are shorter, making them under-represented in the training dataset.
We also report the $R^{2}$ metric, which evaluates how well the predictions fit a regression model given by the optimal diagonal. Precisely, for ground truth $y_{i}$ with mean $\bar{y}$, and predictions $\hat{y}_{i}$
$R^{2}=1-\frac{\Sigma_{i}(y_{i}-\hat{y}_{i})^{2}}{\Sigma_{i}(y_{i}-\bar{y})^{2}}$ (14)
This value ranges from 0 to 1, with 1 being the perfect fit. $\alpha$ and $K$ have a similar score with $R^{2}_{\alpha(t)}=0.75$ and $R^{2}_{K(t)}=0.76$.
In Fig. 10, we plot the training and validation losses as a function of the training epochs, for the network trained on the diffusion parameters. We implement early stopping to prevent overfitting, with a patience of 3 epochs, such that the training stopped in the 7th epoch after no improvements in this time-frame were seen. Training is performed on an HPC system using Tensorflow with CUDA, on an NVIDIA A100-SXM4-40GB GPU, and lasts 164 minutes per epoch on average.
Figure 10: Training and validation loss of the network. Training and validation losses are shown for each epoch in the training procedure of the network that infers $\hat{\alpha}(t)$ and $\hat{K}(t)$. The red circle indicates the epoch the model reverts to when the early stopping is activated. Training loss is scaled by a third for illustration.
Figure 11: Distributions and predictions of the Restricted training subset and the Test set. Predictions vs ground-truth for $\alpha(t)$ [(a) for the restricted training dataset, (b) for the test dataset], and for $K(t)$ [(c) for the restricted training dataset, (d) for the test dataset]. The top panels report histograms of the MAE and MSLE as a function of the ground truth. The density maps report the predicted $\hat{\alpha}(t)$ and $\hat{K}(t)$ vs their respective ground truths. The density is normalised so that each column sums to one. The bottom panels histograms are the distributions of the ground truth parameters of the datasets. Note the logarithmic scaling of $K$.
To investigate the network’s learning, we create an independent test dataset, on a balanced set, featuring an approximately uniform distribution of $\alpha$ with $0<\alpha<2$ and an approximately uniform distribution of $\log K$ for each segment. Trajectories with large $\alpha$ and $K$ rapidly lead the field of view, so that they tend to be observed for a shorter time, leading to the decreasing distribution of $\alpha(t)$ and $K(t)$ shown in the lower panels of Fig. 11. To assess overfitting, we compare the performance to a restricted subset of the training set, with a comparable distribution of $\alpha$ and $K$. Both the restricted training and test datasets have $2\times 10^{4}$ trajectories, with the distributions of point-wise parameters illustrated in the bottom histograms of each subplot in Fig. 11. Note that, despite uniformly sampling the segment parameter distributions, there are some discrepancies between restricted and test sets in the point-wise distributions due to stochastic variations in the bulk experiments the trajectories are created within. Comparing the restricted training and the test data set (left and right density maps in Fig. 11, respectively), we observe comparable performances, which are confirmed by the $\textrm{MAE}_{\alpha(t)}$ and $\textrm{MSLE}_{K(t)}$ metrics reported in Table 1.
Dataset $\textrm{MAE}_{\alpha(t)}$ $\textrm{MSLE}_{K(t)}$ $\textrm{RMSE}_{CP}$ $\textrm{JSC}_{CP}$ $\textrm{MAE}_{\alpha}$ $\textrm{MSLE}_{K}$
Training 0.247 0.140 1.873 0.342 0.274 0.208
Restricted Train 0.185 0.045 1.455 0.705 0.180 0.061
Test 0.191 0.046 1.412 0.737 0.193 0.057
Benchmark 0.191 0.021 1.504 0.577 0.206 0.027
Table 1: Performance of sequandi based on metrics of predictions in different datasets. $\textrm{MAE}_{\alpha(t)}$ and $\textrm{MSLE}_{K(t)}$ are the average of point-wise mean absolute error in $\alpha$ and mean squared logarithmic error in $K$, respectively, of all trajectories in each dataset. $\textrm{RMSE}_{CP}$ is the root mean squared error between true and predicted changepoint times that are true positives (TPs), given in Eq. (9). $\textrm{JSC}_{CP}$ is the rate of these TPs (higher is better), given in Eq. (8). $\textrm{MAE}_{\alpha}$ and $\textrm{MSLE}_{K}$ are the average errors of $\alpha$ and $K$ only in the ensemble of TP segment detections.
These metrics (closely related to the ones used during training) show how the restricted training and the test datasets have similar performances, demonstrating that the training did not lead to overfitting. For completeness, we also report in Table 1 the metrics used in the AnDi challenge segment-wise inference, using the paired segments as described in Sections 4.3 and  5.2. These evaluate predictions from the post-processed pooling of $\hat{\alpha}(t),\hat{K}(t)$ with the $\hat{t}^{\textrm{CP}}$ trained on a different network. These metrics are interesting for evaluating how the networks perform on different datasets but do not provide immediate information about overfitting.
A.1 Computation times of predictions
While training is a time-consuming process, using the trained networks to make predictions is relatively fast. To infer $\hat{\alpha}(t)$ and $\hat{K}(t)$, we must use the networks three times per trajectory set, as described in Section 3.2; once for the original trajectory, and twice more for a shift of one and two increments. The three predictions (with three time-step resolution) are pooled to obtain the refined $\hat{\alpha}(t)$ and $\hat{K}(t)$ traces with a single time-step resolution. Using the sequandi networks on a laptop (13th Gen Intel Core i7 processor, 32 GB RAM), it takes 8 minutes for the test dataset predictions pictured in Fig. 11(b) to be made. Comparatively, on this dataset it takes 20 minutes to find all $\hat{\alpha}(t)$ and $\hat{K}(t)$ using the sliding window TA-MSD technique with linear fits.
Appendix B Hyperparameter fine-tuning
We chose our architecture combining information from the literature, [11] and trial and error. In this section, we compare our performance with the one of a network where we systematically perform a hyperparameter search with the hyperband method [102] in Keras. This is a tuning method which searches randomly at first, trying different networks for short amounts of time, before going deeper in the training with the best candidates. The candidates are whittled down successively like in a sports bracket, until the best one remains. We perform the search both on the network designed for diffusion parameter inference and for the one for changepoint locations’ prediction, with 90 combinations of hyperparameters being tested in each case. The best architecture for the $\alpha(t)$ and $K(t)$ inference is: [LSTM Dimensions 320 and 160, recurrent dropout 0.2, dropout 0.1, learning rate 0.003]. For the $t^{\textrm{CP}}$ inference: [LSTM Dimensions 380 and 150, recurrent dropout 0.1, dropout 0.2, learning rate 0.0005]. These architectures are not too different from the one of sequandi. We train them on the full training set, and find that their performance is comparable to that of sequandi, as reported in Table 2 on the four datasets. Interestingly, sequandi performs better on the experimentally-informed benchmark dataset (compare with Table 1), which is the one where we are interested in optimal performance.
Dataset $\textrm{MAE}_{\alpha(t)}$ $\textrm{MSLE}_{K(t)}$ $\textrm{RMSE}_{CP}$ $\textrm{JSC}_{CP}$ $\textrm{MAE}_{\alpha}$ $\textrm{MSLE}_{K}$
Training 0.228 0.122 1.947 0.293 0.258 0.191
Restricted Train 0.188 0.063 1.617 0.684 0.177 0.076
Test 0.194 0.064 1.421 0.699 0.187 0.075
Benchmark 0.208 0.028 1.509 0.483 0.217 0.031
Table 2: Performance of the tuned networks based on metrics of predictions in different datasets. $\textrm{MAE}_{\alpha(t)}$ and $\textrm{MSLE}_{K(t)}$ are the average of point-wise mean absolute error in $\alpha$ and mean squared logarithmic error in $K$, respectively, of all trajectories in each dataset. From the AnDi challenge, $\textrm{RMSE}_{CP}$ is the root mean squared error between true and predicted changepoint times that are true positives (TPs). $\textrm{JSC}_{CP}$ is the rate of these TPs (higher is better). $\textrm{MAE}_{\alpha}$ and $\textrm{MSLE}_{K}$ are the average errors of $\alpha$ and $K$ only in the ensemble of TP segment detections.
Appendix C Neural network changepoint detection signal
@@
Appendix D Changepoint algorithm optimisation
@@
    The first step checks if there are changepoints in a trajectory segment. Here, we are using as input the prediction from our networks $\hat{\alpha}(t)$. A straight line $\alpha_{\text{fit}}(t)$ is fitted. The residual is computed as $d(t)=\hat{\alpha}(t)-\alpha_{\text{fit}}(t)$, followed by the cumulative sum $\text{CUMSUM}(t)$ of $d(t)$. The statistic $D_{\text{data}}$ is the difference between the maximum and minimum values of $\text{CUMSUM}(t)$. The significance of $D_{\text{data}}$ is tested by permuting $d(t)$ multiple times to generate a distribution of $\text{CUMSUM}(t)$ curves and $D_{\text{permute}}$ values. A confidence level (e.g., 95%) sets the threshold $D_{\text{cutoff}}$. If $D_{\text{data}}\geq D_{\text{cutoff}}$, a changepoint is present. We discuss the relevance of the confidence level in D.1
@@
D.1 Confidence levels for the CPDA